1.    A lubricated shaft of 200mm radius rotates inside a concentric sleeve bearing of 205mm radius. Determine the viscosity of the fluid, which fills the space between the sleeve and the shaft if a torque of 1.5N.m is required to maintain an angular velocity of 120rpm.take length of sleeve as 450mm.



1
Expert's answer
2021-08-16T17:31:06-0400

Gives

∆r=205mm-200mm=5mm

Torque=1.5Nm

Hight(h)=450mm

"w=2\pi\times\frac{120}{160}=4\pi rad/sec"

Linear velocity

"v=rw"

"\frac{dv}{dr}=\frac{v}{r}=\frac{rw}{r}=w" =4"\pi"

"\tau=rF"

"F=\frac{\tau}{r}"

"F=\frac{1.5}{205\times10^{-3}}=7.31N"

Area="2\pi rh"

Put value


"A=2\pi \times205\times10^{-3}450\times10^{-3}=0.58m^2"

Newton's law

Force

"F=-\eta A\frac{dv}{dr}"

"\eta=\frac{F}{A\frac{dv}{dr}}"

"\eta=\frac{7.31}{{0.58}\times4\times3.14}=1.0034" "Nsec\/m^2"